Ed invested twice as much in a 4% bond as he invested in a 3% bond. His total yearly interest from the two investments was $770. How much did he invest in each bond?
Answer by checkley77(12844)

.04*2x+.03x=770
.08x+.03x=770
.11x=770
x=770/.11
x=7,000 invested @ 3%
2*7,000=14,000 invested @ 4%
Proof:
.04*14,000+.03*7,000=770
560+210=770
770=770